Bats use an ultrasonic sound to locate their prey in the dark. Its brief squeak starts at around 25,000 Hz and slips an octave up to 50,000 Hz. Show that the wavelengths in this signal are suitable for locating insects measuring 6 to 12 millimeters.
Answer #1

minimum wavelength

L1 = v / f1 = 340 / 50000 = 6 mm

max wavelength

L2 = v / f2 = 340 / 25000 = 12 mm

so the objects which are comparable to the wavelength of the wave can be recognised.

=======

Comment before rate in case any doubt, will reply for sure.. goodluck